the organic garden at a restaurant is rectangular in shape and has a perimeter of 84 feet, if the length of the garden is 24 feet longer than the width what is the area of the garden in square feet
Answer by TimothyLamb(4379)

---
x = length
y = width
---
2x + 2y = 84
x = y + 24
---
put the system of linear equations into standard form
---
2x + 2y = 84
x - y = 24
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = length = 33 ft
y = width = 9 ft
area = xy
area = 33*9
area = 297 sq.ft
---
system matrix
2 2
1 -1
---
inverse of system matrix
0.25 0.5
0.25 -0.5
---
determinant of system matrix = -4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php